3. Use Euclids algorithm to compute the following. Show all your steps 1. gcd(781, 994) 2. gcd(67457, 43521)
Answer #1

Euclids Algorithm page-( Recall some Basic fact - for any positive Integer a and b good (a, b) ffa >b a= bqtr orchi b = 192+page- fos g.c.d (701,994) 994 = (101(1) + 213 781 = (213) (3) + 142 213 = (142)(1 + 71 142. = 671) (2) + 0 O< 213 < 781 0 < 1Page- quica (67 457, 43521) 67457 = (43521 (0 + 23936 43521 = (2.3936) (1) + 19505 23936 = (19505) ( + 435| 19585 = (4351) (4page- For more Explainatia 43521) 6745711 43521 23936) 4352111 23.936 1.9505)2393611 19505 4351) 1958 5(4 17404 210143516 21